the square of a negative number is 48 more than twice the negative number. find the number
Answer by rwm(914)

x^2=2x+48
x^2-2x-48=0
(x-8)(x+6)=0
x=8 x=-6
x must be negative so x=-6
check
36=48+-12
36=36